The attractive force between two equal but opposite charges is F newtons when the charges are 6 cm apart. What should be their distance apart to increase the attractive force to 6F newtons?

1
Expert's answer
2022-03-21T12:44:50-0400

Attractive force

"F=\frac{kq_1q_2}{r^2}"

"F=-\frac{kqq}{r^2}"

"\frac{F_1}{F_2}=\frac{r_2^2}{r_1^2}"

"\frac{F}{6F}=\frac{r_2^2}{0.06^2}"

"r_2^2=\frac{3.6\times10^{-3}}{6}"

"r=0.0244m"

"r=2.44cm"